OpenStudy (sloppycanada):

A ladder is leaning up against a 14-foot wall at an angle of elevation of 68°. How far is the foot of the ladder from the wall? Round your answer to the nearest tenth of a foot.

OpenStudy (sloppycanada):



OpenStudy (anonymous):

 do you remember your trig relations

OpenStudy (sloppycanada):

I should be using Tangent yeah? So Opposite over Adjacent. 14/x tan14 x 68?

OpenStudy (sloppycanada):

wait sorry tan 68 x 14.

OpenStudy (anonymous):

\(\tan68=\dfrac{14}{x}\\2.47508=\dfrac{14}{x}\\x=\dfrac{14}{2.47508}\)

OpenStudy (sloppycanada):

I don't understand why I always get that mixed up. If x was on the top, then it would be multiplied.. x = 5.656 x = 5.7

OpenStudy (anonymous):

remember that the relation is revolving around the angle we're working with. so tan of 68 is opp/adj looks good